OpenStudy (anonymous):

The table shows the annual consumption of cheese per person in the United States for selected years in the 20th century. Let x = number of years since 1900, and y = pounds per person. What cubic model best fits this data? Year Pounds Consumed 1902 1.959 1924 6.373 1946 7.29 1982 15.817 A. y = 0.00989x3 - 0.0000872x2 + 1.192x + 0.403 B. y = -0.0000872x3 + 0.00989x2 - 0.403x - 1.192 C. y = -0.0000872x3 - 0.00989x2 - 0.403x + 1.192 D. y = 0.0000872x3 - 0.00989x2 + 0.403x + 1.192

OpenStudy (anonymous):

can you help me.?

OpenStudy (richyw):

yes just a second.

OpenStudy (anonymous):

okayy thanks

OpenStudy (richyw):

ok. so right now you have a table that shows the year. but what you want is to rewrite that table with the years since 1900.

OpenStudy (anonymous):

okay how would i do that can you show me

OpenStudy (richyw):

then after that, try plugging those numbers into each equation.

OpenStudy (richyw):

if the year is 1902, then how many years have passed since 1900?

OpenStudy (anonymous):

2

OpenStudy (richyw):

so for the year 1902 you have x=2

OpenStudy (anonymous):

okay so do i plug that into something.?

OpenStudy (richyw):

you can just plug 2 into each of those equations and see what they give you. That's a time consuming way to do it, but the easiest.

OpenStudy (anonymous):

so i would pulg it inot this 1st right A. y = 0.00989x3 - 0.0000872x2 + 1.192x + 0.403

OpenStudy (anonymous):

into*

OpenStudy (richyw):

yup. see what each one gives you.

OpenStudy (anonymous):

okay so i just plug it in where i see x right..? im sorry i have no math skills lol. but im trying

OpenStudy (richyw):

yes, x = 2 so you plug it in where you see x.

OpenStudy (anonymous):

im sorry but that doest help me figure it out it just adds a 2 to everything :/

OpenStudy (anonymous):

A. y = 0.0098923 - 0.000087222 + 1.1922 + 0.403

OpenStudy (anonymous):

1.605005078 thats what i got :x

OpenStudy (richyw):

i was assuming that the you copied and pasted this and the first function was\[y = 0.00989x^3 - 0.0000872x^2 + 1.192x + 0.403\]

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

http://gyazo.com/3b5078294c16e643ba5aebea38b18ff6

OpenStudy (anonymous):

thats the entire problem

OpenStudy (richyw):

yeah so that means you have \[0.00989\times 2^3 -0.0000872\times2^2+ \dots\]

OpenStudy (anonymous):

okay so i would just multiply from this point

OpenStudy (richyw):

yes. you plug it in. what you are trying to do is see which equation matches the data in the table.

OpenStudy (richyw):

the blue dots in this graph is the data provided in the table, and the four functions are plotted. notice how one fits the data nicely? your goal is to figure out which one does. The easiest way to do this is to just check each formula at one point, in this case x=2.

OpenStudy (richyw):

If you practice enough math, you can figure out quicker ways to do this, but they are more difficult than just plugging in...

OpenStudy (anonymous):

okay so im guessing that from looking at the graph that the answer is pink line

OpenStudy (anonymous):

which would be answer C

OpenStudy (richyw):

wait what? why the pink line? and I did not tell you which line corresponded to which function. I was just trying to help you understand what you are doing.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i was guessing

OpenStudy (richyw):

here. the first one would look like this \[0.00989\times 2^3 - 0.0000872 \times 2^2 + 1.192\times 2 + 0.403=2.866\]

OpenStudy (richyw):

does that match the data given to you?

OpenStudy (anonymous):

no

OpenStudy (richyw):

ok. so try doing that with the next equation on the list of answers. see if it matches your data

OpenStudy (anonymous):

it doesnt

OpenStudy (richyw):

what number did you get?

OpenStudy (anonymous):

186

OpenStudy (richyw):

you did something wrong. you know the order of operations right?

OpenStudy (anonymous):

yes

OpenStudy (richyw):

check your calculations again then.